Hello Again!
Can someone help me solve this problem?
\/2x-1-x=-2
(\/ is the radical sign) the radical sign is only over 2x-1
Thanks in advance for your help!
ac

Answer by asha(30)

given square root of 2x-1=x-2
squaring both sides we get
2x-1 = (x-2)^2
2x-1 = x^2 -4x +4 taking all the terms to one side
we get x^2-4x-2x+4+1=0
x^2 -6x +5=0
splitting the middle term so that sum is -6 and product is 5 we get
x^2-5x-x+5=0
x(x-5)-1(x-5)=0
(x-1)(x-5)=0
x=1 or x=5 is the answer
good luck!!!